OpenStudy (anonymous):

Polar coordinates of a point are given. Find the rectangular coordinates of the point. (2.1 ; 2pi/9) A. (-1.6, -1.3) B. (1.6, 1.3) C. (1.3, 1.6) D. (-1.3, -1.6)

OpenStudy (anonymous):

It is B. look on this website http://www.wolframalpha.com/input/?i=x%3D2.1 \cos%28\frac{2\pi}{9}%29%2Cy%3D2.1\sin%28\frac{2\pi}{9}%29

OpenStudy (anonymous):

OK THANKS ;-) @bloodrain

OpenStudy (anonymous):

yw, any time :D